OpenStudy (anonymous):

(square root of 9v-15 = square root of 7v-5.) there could be multiple solutions or no solutions PLEASE HELP

OpenStudy (anonymous):

\[\sqrt{9v-15}=\sqrt{7v-5}~~?\]

OpenStudy (anonymous):

yes! :)

OpenStudy (anonymous):

Square both sides and you get \[9v-15=7v-5~~\Rightarrow~~2v=10~~\Rightarrow~~v=5\] Before you say this is a solution, check to make sure it's not extraneous: \[\sqrt{9\times5-15}=\sqrt{45-15}=\sqrt{30}\] \[\sqrt{7\times5-5}=\sqrt{35-5}=\sqrt{30}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw